Flower color is determined by a single locus, where the heterozygote displays an intermediate phenotype. Using the numbers below observed for the 3 phenotypes, calculate p and q Is this population likely to be in H-W equilibrium?

Number of individuals Phenotypes Red 200 50 Pink White 750
Answer #1

Answer:

q = 0.775

p = 0.225

Explanation:

Red = RR = 200

Pink = Rr = 50

White = rr = 750

RR 200 = 200*2 = 400 “R” alleles

Rr 50 = 50 R & 50 r alleles

rr 750 = 1500 “r” alleles

Total allele= 2000

Total R alleles = 400+50 = 450

Frequency of R allele = = p = 450/2000 = 0.225

Total r alleles = 1500+50 = 1550

Frequency of r allele =q= 1550/2000 = 0.775